OpenStudy (anonymous):

A line passes through the point (-3, 2) and has a slope of 1/3. Which of the following points also lies on this line

OpenStudy (anonymous):

Ok, will you please post the points?

OpenStudy (anonymous):

(-2, 5) (-9, 1) (5, 6) (6, 5)

OpenStudy (anonymous):

So, to solve this, you have to figure out the equation of your line

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

The equation of a line is y=mx+b, this may seem difficult, however, you have to know that m means slope and b means y intercept

OpenStudy (anonymous):

The problem gives you the slope, 1/3. So we "plug it in". So now, we have y=1/3 x+b. To solve find the value of b to complete the line equation, we are going to use the point they gave us,(-3,2).

OpenStudy (anonymous):

b=2?

OpenStudy (anonymous):

No, y=1/3x+b. 2=-1+b, b=3. Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so the equation of our line is y=1/3 x+3. Now, we will "plug in the points" to try to find our answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lets start with (-2,5) 5=-2/3+3. Is this true?